According to the World Bank, China’s gross national income (GNI) in 2012 was $7,748.9 billion while its PPP-adjusted (or PPP-implied) GNI was $12,435.4 billion (World Bank, World Development Report, 2014). Based on these numbers, what was the real exchange rate of the Chinese currency?
Answer #1

The solution to the given question is provided below:

Real exchange rate=GNI/GNI PPP adjusted=7748.9/12435.4=0.6231